# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ II-C Post-training and Alignment Strategies
Kumar et al. [33] delves into post-training techniques for LLMs, emphasizing that while pre-training establishes a robust linguistic foundation, it is subsequent refinement - through methods such as fine-tuning, reinforcement learning, and testing-time scaling - that truly enhances the models’ reasoning abilities, factual accuracy, and overall adaptability. The work systematically examines these post-training strategies, focusing on their potential to address critical challenges such as catastrophic forgetting, reward hacking, and the trade-offs encountered during inference. Furthermore, it highlights emerging avenues in model alignment and scalable adaptation, underscoring the importance of these techniques in ensuring that LLMs understand language and perform effectively across a diverse range of real-world tasks.
The survey by Wang et al. [32] offers a systematic overview of various alignment techniques developed to enhance the reliability of LLMs. It emphasizes that despite LLMs achieving remarkable capabilities through massive pretraining, the variability in data quality can still lead to suboptimal or undesired outputs. To mitigate this, the paper categorizes and examines a range of methods—including reinforcement learning from human feedback (RLHF), reinforcement learning from AI feedback (RLAIF), proximal policy optimization (PPO), and direct preference optimization (DPO)—detailing how each approach contributes to aligning model outputs with human expectations. By organizing these techniques into distinct topics, the survey fills a critical gap in the literature and provides a comprehensive framework that not only aids in understanding current alignment strategies but also guides future research toward improving the performance and safety of LLMs.
The survey by Tie et al. [34] presents a thorough synthesis of post-training methodologies to overcome the inherent limitations of pre-trained large language models. It critically addresses challenges such as constrained reasoning abilities, ethical ambiguities, and performance issues in specialized domains, underscoring the need for advanced post-training strategies. The authors systematically categorize the evolution of post-training language models into five core paradigms: fine-tuning to enhance task-specific accuracy, alignment to better meet human expectations, reasoning to facilitate multi-step inference despite reward design challenges, efficiency to optimize resource use amid growing model complexity, and integration and adaptation to extend capabilities across diverse modalities. By mapping progress from early alignment techniques exemplified by ChatGPT to the innovative reasoning approaches of models like DeepSeek-R1, the survey establishes a structured taxonomy that not only consolidates recent advancements but also charts a strategic path for future research toward developing LLMs with improved precision, ethical robustness, and domain adaptability.
@@ II-D Advanced Reasoning and Chain-of-Thought
The survey by plaat et al. [28] examines the evolving landscape of reasoning with LLMs as they scale to billions of parameters, enabling advanced in-context learning and few-shot capabilities. It highlights the breakthrough performance of these models on traditional ”System 1” tasks—such as translation, summarization, and question-answering—and extends the discussion to ”System 2” reasoning enabled by chain-of-thought prompting. The authors present a comprehensive taxonomy that categorizes various strategies for generating, evaluating, and controlling multi-step reasoning processes. In doing so, the paper offers an in-depth review of current methodologies and identifies key open challenges, including integrating sequential decision processes and reinforcement learning. Furthermore, it explores the potential for self-improvement and metacognitive abilities through prompt-based techniques while noting that LLMs’ ultimate shift toward autonomous reasoning remains a promising avenue for future research.
Chen et al. [36] provide a unified perspective on the emerging paradigm of long chain-of-thought (Long CoT) reasoning within LLMs. Building on the impressive capabilities demonstrated by models such as OpenAI-O1 and DeepSeek-R1 in complex domains like mathematics and coding, the authors distinguish Long CoT from traditional short chain-of-thought approaches by introducing a novel taxonomy of reasoning paradigms. The paper highlights key characteristics of Long CoT—deep reasoning, extensive exploration, and reflective processes—that enable models to solve more intricate tasks and yield more coherent outcomes. It also examines overthinking and test-time scaling phenomena, offering insights into how these behaviors manifest in practice. By identifying significant research gaps and proposing future directions, including integrating multi-modal reasoning and efficiency improvements, the survey sets a strategic roadmap for advancing logical reasoning capabilities in artificial intelligence.
@@ II-E Evaluation of LLM-based Agents
Yehudai et al. [37] provide a comprehensive review of evaluation methodologies for LLM-based agents, marking a significant paradigm shift in autonomous AI systems. These agents, which are capable of planning, reasoning, tool usage, and maintaining memory in dynamic environments, require robust evaluation frameworks to gauge their performance accurately. The paper systematically dissects evaluation benchmarks and methodologies across four critical dimensions: assessing fundamental capabilities such as planning, self-reflection, tool utilization, and memory; domain-specific benchmarks for areas like web interactions, software engineering, scientific inquiry, and conversational tasks; evaluation standards for generalist agents; and overarching frameworks that integrate these diverse facets. Notably, the survey highlights emerging trends toward more realistic, challenging evaluations with continuously updated benchmarks while identifying key gaps—especially regarding cost-efficiency, safety, robustness, and scalability—that must be addressed. This work thus not only maps the rapidly evolving landscape of agent evaluation but sets forth a clear research agenda for advancing future methodologies.
@@ II-F Our Survey
Our study makes a series of novel contributions compared to related works as presented in Tab. I. We analyze the top 27 LLM models released between 2023 and 2025—including models such as Mistral AI Small 3 24B, DeepSeek-R1, Search-o1, QwQ-32B, and phi-4—providing detailed insights into their architectures, performance, and innovations. In addition, we offer a thorough overview of training methodologies by categorizing approaches into general training techniques, mixture-of-experts (MoE) innovations, retrieval-augmented generation (RAG), chain-of-thought and self-improvement methods, as well as test-time compute scaling, distillation, and reinforcement learning (RL) strategies.
@@
Model Name Year Parameters Key Metrics/Benchmarks Innovations/Techniques Observations UI-TARS [39] 2025 2B, 7B, 72B OSWorld: 24.6 (50-step), 22.7 (15-step); AndroidWorld: 46.6 (vs. GPT-4’s 34.5) Native GUI agent using screenshot inputs; end-to-end design; System-2 reasoning; unified action modeling; iterative training GUI agent outperforming comparable models (e.g., Claude) Tülu 3 405B [40] 2025 405B Surpasses Llama 3.1-Instruct, Qwen 2.5, Mistral, GPT-4o-mini, Claude 3.5-Haiku Transparent, reproducible post-training using SFT, DPO, and RLVR Built on Llama 3.1 base; open-source pipeline Mistral AI Small 3 24B [41] 2025 24B 81% accuracy on MMLU; 150 tokens/s; 3$\times$ faster on same hardware Fewer layers for lower latency; optimized for instruction-following; no RL or synthetic data Efficient LLM for generative AI tasks DeepSeek-R1 [42] 2025 1.5B–70B FP8 quantization cuts memory by 75%; MTP yields 2–3$\times$ faster token generation Uses Mixture of Experts (MoE), Multihead Latent Attention (MLA), and Multi-Token Prediction (MTP) Matches OpenAI’s o1-1217 in reasoning tasks Qwen 2.5 VL [43] 2024 3B, 7B, 72B 72B version matches GPT-4o and Claude 3.5-Sonnet on multimodal benchmarks Naive Dynamic Resolution (NDR) for adaptive visual tokenization; M-RoPE; unified image–video processing Advanced vision–language integration HuatuoGPT-o1 [20] 2024 70B $\sim$80improvement on MedQA and PubMedQA; complex reasoning chains $>$700 tokens Two-stage training: 40K verifiable problems then PPO-based RL with a robust verifier Focused on enhancing medical reasoning Search-o1 [5] 2025 Outperforms base models by 4.7% and traditional RAG systems by up to 29.6% on multi-hop questions Retrieval-augmented generation with a “Reason-in-Documents” module Enhances long stepwise reasoning in LRMs InternLM2 [44] 2024 8B (up to 128k context) 83.0% on MATH-500; 20.0% on AIME2024; trained on 4T tokens COOL RLHF; multi-phase pre-training (4k $\rightarrow$ 32k tokens) for long-context learning Open-source LLM with 75% cost savings MiniCPM-o 2.6 [45] 2024 8B Visual: 70.2 on OpenCompass; encodes 1.8M pixels in 640 tokens On-device multimodal processing; real-time speech and OCR Designed for on-device applications with flexible voice customization KaLM-Embedding [46] 2025 $<$1B Average score: 64.53 on MTEB benchmark Persona-based synthetic data generation; ranking consistency filtering; Matryoshka Representation Learning Multilingual embedding model with flexible dimensions rStar-Math [19] 2025 7B 90.0% on MATH; 53.3% on AIME Monte Carlo Tree Search with code-augmented chain-of-thought; self-evolution; improved PRM training Specialized for advanced mathematical reasoning phi-4 [47] 2024 14B Trained on 9.8T tokens; 16K token context; comparable to 70B models High-quality synthetic data via multi-agent self-revision; SFT and DPO for post-training STEM-focused QA model (Microsoft) DeepSeek-V3 [48] 2024 671B total (37B active) Throughput: 60 tokens/s; 2.788M H800 GPU hrs (\$5.5M) MoE, MLA, DeepSeekMoE; auxiliary-loss-free load balancing; multi-token prediction; FP8 training Massive LLM for advanced reasoning DeepSeekMath 7B [49] 2024 7B 51.7% on MATH (60.9% with self-consistency) Group Relative Policy Optimization (GRPO); tailored data selection pipeline Specialized for mathematical reasoning Qwen 2.5 [50] 2024 0.5B–72B 72B-Instruct outperforms many larger models on multiple benchmarks Enhanced pre-training (7T $\rightarrow$ 18T tokens), extensive SFT, offline RL (DPO) and online RL (GRPO); modern features (GQA, SwiGLU, RoPE) Comprehensive general-purpose LLM series
@@
Model Name Year Parameters Key Metrics/Benchmarks Innovations/Techniques Observations ModernBERT [51] 2024 139M (base), 395M (large) Extended context: 8,192 tokens; trained on 2T tokens; 2–4$\times$ faster than peers Flash Attention 2; RoPE embeddings; alternating attention mechanism Next-generation encoder for classification and retrieval Gemini 2.0 Flash [52] 2024 51.8% on SWE-bench Verified Multimodal Live API for real-time audio/video streaming; native tool use; supports 8 voices and 109 languages Developed by Google DeepMind for enhanced developer applications Llama 3.3 70B-Instruct [53] 2023 70B $\sim$4overall improvement; +9% on MATH; +8% on HumanEval Improved post-training techniques without altering the architecture Matches performance of high-end models (e.g., GPT-4o, Anthropic) Llama 3.1 [54] 2024 8B, 70B, 405B 405B variant matches/exceeds GPT-4o benchmarks; supports 128k token context Trained on $>$15T tokens with 25M human/synthetic samples; optimized quantization (FP8, AWQ, GPTQ) Offered in Instruct and Base versions; commercial-friendly Amazon Nova [55] 2023 Multiple variants Context up to 300k tokens; $>$200 languages; Pricing: Micro (\$0.035 input, \$0.14 output), Lite (\$0.06/\$0.24), Pro (\$0.80/\$3.20) Suite includes text-only, multimodal, and specialized models (Canvas for image, Reel for video); watermarking Deployed on AWS (US); integrates with Amazon Bedrock QwQ-32B-Preview [56] 2023 32.5B; 32,768 token context GPQA: 65.2%; AIME: 50.0%; MATH-500: 90.6%; LiveCodeBench: 50.0% Open-weight reasoning model based on Qwen2.5; experimental with noted limitations Released under Apache 2.0 on Hugging Face Qwen2.5-Coder-32B [57] 2023 32B; 128K token context HumanEval: 92.7 (vs. 92.1); EvalPlus: 86.3 (vs. 85.9); Aider (code repair): 73.7 Code-focused architecture with extended system prompts; multilingual support ($>$40 languages) Optimized for diverse coding benchmarks Hunyuan-Large [58] 2024 389B total (52B active); up to 256K tokens Uses 1.5T synthetic tokens (of 7T total) Largest open-source MoE with 16 experts; mixed expert routing; KV cache compression; expert-specific learning rates Outperforms Llama 3.1-70B; competitive with 405B models Granite 3.0 [59] 2023 8B & 2B Dense; 3B & 1B MoE OpenLLM Leaderboard: 37.6 (vs. Llama 3.1 8B: 37.3) Speculator Head accelerates token processing up to 220%; multilingual, coding, function-calling features Lightweight foundation for enterprise/on-device; Apache 2.0 licensed Pyramid Flow SD3 [60] 2024 2B (DiT model) Videos up to 10 sec at 768p, 24 FPS; Training: 20.7k A100 GPU hrs Unified pyramidal flow matching; autoregressive temporal pyramid for spatiotemporal compression First robust open-source text-to-video and image-to-video model Molmo model [61] 2024 7B MoE (A1B), dual 7B, 72B Molmo 72B outperforms Llama 3.2 90B and Pixtral 12B; matches proprietary models (GPT-4o, Gemini 1.5 Pro, Claude Sonnet 3.5) Built entirely from scratch using the PixMo dataset; uses OpenAI CLIP in the 72B variant Open-weight vision-language model family Moshi [62] 2024 7B Theoretical latency: 160 ms (200 ms on L4 GPU); 12.5 Hz at 1.1 kbps Full-duplex spoken dialogue via a Temporal Transformer; “Inner Monologue” for time-aligned text prefix to audio tokens Speech–text model for real-time, on-device dialogue (cc-by-4.0)
@@ III Top LLM models
Recent advances in LLMs during 2023–2025 have demonstrated remarkable progress in architecture and training methodologies. For example, UI-TARS [39] and Tülu 3 [40] exemplify novel approaches in multimodal processing, and transparent post-training pipelines, respectively, that have enabled significant improvements in reasoning, code generation, and domain-specific tasks. These developments leverage innovations such as reinforcement learning with verifiable rewards, retrieval-augmented generation, and large-scale fine-tuning, thereby addressing critical scalability, efficiency, and task adaptability challenges. The concentrated efforts during this period underscore a broader trend toward enhancing performance through high-quality synthetic data and advanced optimization techniques.
In parallel, models such as HuatuoGPT-o1 [20] and Search-o1 [5] highlight a focused commitment to refining reasoning capabilities in specialized domains, including medicine and long-step problem-solving. In addition, breakthroughs in vision language integration and context management have been achieved by systems such as Qwen2.5 VL [43] and InternLM2 [44], further broadening the scope of applications for LLM. These contributions, emerging in a concentrated timeframe between 2024 and 2025, reinforce the rapid pace of innovation in the field and provide a robust foundation for subsequent research and development in artificial intelligence.
@@ III-A UI-TARS model
UI-TARS [39] is a novel native GUI agent model that exclusively utilizes screenshot inputs to execute human-like interactions such as mouse and keyboard operations. Unlike conventional frameworks that rely on heavily wrapped commercial models and meticulously designed prompts, UI-TARS has been developed as an end-to-end system that surpasses these sophisticated performance approaches. Experimental evaluations across more than ten GUI agent benchmarks—assessing perception, grounding, and task execution—demonstrate its state-of-the-art performance. For example, on the OSWorld benchmark, UI-TARS achieves scores of 24.6 and 22.7 for 50-step and 15-step evaluations, respectively, outperforming comparable models like Claude, while on AndroidWorld, it scores 46.6 compared to GPT-4’s 34.5. The system integrates several key innovations: it employs enhanced perception through a large-scale dataset of GUI screenshots for context-aware understanding; it utilizes unified action modeling to standardize interactions across platforms; it incorporates System-2 reasoning to enable deliberate multi-step decision-making via task decomposition, reflection, and milestone recognition; and it leverages iterative training with reflective online traces to continuously refine its performance with minimal human intervention. Available in 2B, 7B, and 72B parameter configurations, UI-TARS exemplifies a significant step forward in building versatile, reasoning-capable GUI agents that can perform complex interactions across diverse computing environments.
@@ III-B Tülu 3 405B model
The recent study on Tulu 3 [40] presents a significant advancement in language model post-training, offering a fully transparent and reproducible framework for refining and enhancing LLMs. Unlike proprietary post-training pipelines, which often lack transparency regarding training data and methodologies, Tulu 3 provides a comprehensive open-source alternative, sharing the model weights and the complete post-training recipe, including datasets, training code, infrastructure, and evaluation tools. Built upon Llama 3.1 base models, Tulu 3 surpasses existing instruct-tuned models such as Llama 3.1-Instruct, Qwen 2.5, Mistral, and even proprietary models like GPT-4o-mini and Claude 3.5-Haiku in multiple benchmarks. The training methodology incorporates Supervised Fine-tuning (SFT), Direct Preference Optimization (DPO), and an innovative technique termed Reinforcement Learning with Verifiable Rewards (RLVR), which ensures reliable and measurable improvements in model behavior. To further enhance post-training effectiveness, the authors introduce a multi-task evaluation framework, which includes development and unseen benchmarks, standardized performance metrics, and rigorous decontamination of existing open datasets.

### v2

@@ II-C Post-training and Alignment Strategies
Kumar et al. [kumar2025llm] delves into post-training techniques for LLMs, emphasizing that while pre-training establishes a robust linguistic foundation, it is subsequent refinement, through methods such as fine-tuning, reinforcement learning, and testing-time scaling, that truly enhances the models’ reasoning abilities, factual accuracy, and overall adaptability. The work systematically examines these post-training strategies, focusing on their potential to address critical challenges such as catastrophic forgetting, reward hacking, and the trade-offs encountered during inference. Furthermore, it highlights emerging avenues in model alignment and scalable adaptation, underscoring the importance of these techniques in ensuring that LLMs understand language and perform effectively across a diverse range of real-world tasks.
The survey by Wang et al. [wang2024comprehensive] offers a systematic overview of various alignment techniques developed to enhance the reliability of LLMs. It emphasizes that despite LLMs achieving remarkable capabilities through massive pretraining, the variability in data quality can still lead to suboptimal or undesired outputs. To mitigate this, the paper categorizes and examines a range of methods, including reinforcement learning from human feedback (RLHF), reinforcement learning from AI feedback (RLAIF), proximal policy optimization (PPO), and direct preference optimization (DPO), detailing how each approach contributes to aligning model outputs with human expectations. The survey fills a critical gap in the literature by organizing these techniques into distinct topics. It provides a comprehensive framework that not only aids in understanding current alignment strategies but also guides future research toward improving the performance and safety of LLMs.
The survey by Tie et al. [tie2025survey] presents a thorough synthesis of post-training methodologies to overcome the inherent limitations of pre-trained large language models. It critically addresses challenges such as constrained reasoning abilities, ethical ambiguities, and performance issues in specialized domains, underscoring the need for advanced post-training strategies. The authors systematically categorize the evolution of post-training language models into five core paradigms: fine-tuning to enhance task-specific accuracy, alignment to better meet human expectations, reasoning to facilitate multi-step inference despite reward design challenges, efficiency to optimize resource use amid growing model complexity, and integration and adaptation to extend capabilities across diverse modalities. By mapping progress from early alignment techniques exemplified by ChatGPT to the innovative reasoning approaches of models like DeepSeek-R1, the survey establishes a structured taxonomy that not only consolidates recent advancements but also charts a strategic path for future research toward developing LLMs with improved precision, ethical robustness, and domain adaptability.
@@ II-D Advanced Reasoning and Chain-of-Thought
The survey by Plaat et al. [plaat2024reasoning] examines the evolving reasoning landscape with LLMs as they scale to billions of parameters, enabling advanced in-context learning and few-shot capabilities. It highlights the breakthrough performance of these models on traditional ”System 1” tasks such as translation, summarization, and question-answering. It extends the discussion to ”System 2” reasoning enabled by chain-of-thought prompting. The authors present a comprehensive taxonomy that categorizes various strategies for generating, evaluating, and controlling multi-step reasoning processes. In doing so, the paper offers an in-depth review of current methodologies and identifies key open challenges, including integrating sequential decision processes and reinforcement learning. Furthermore, it explores the potential for self-improvement and metacognitive abilities through prompt-based techniques while noting that LLMs’ ultimate shift toward autonomous reasoning remains a promising avenue for future research.
Chen et al. [chen2025towards] provide a unified perspective on the emerging paradigm of long chain-of-thought (Long CoT) reasoning within LLMs. Building on the impressive capabilities demonstrated by models such as OpenAI-O1 and DeepSeek-R1 in complex domains like mathematics and coding, the authors distinguish Long CoT from traditional short chain-of-thought approaches by introducing a novel taxonomy of reasoning paradigms. The paper highlights key characteristics of Long CoT deep reasoning, extensive exploration, and reflective processes that enable models to solve more intricate tasks and yield more coherent outcomes. It also examines overthinking and test-time scaling phenomena, offering insights into how these behaviors manifest in practice. By identifying significant research gaps and proposing future directions, including integrating multi-modal reasoning and efficiency improvements, the survey sets a strategic roadmap for advancing logical reasoning capabilities in artificial intelligence.
@@ II-E Evaluation of LLM-based Agents
Yehudai et al. [yehudai2025surveyevaluationllmbasedagents] provide a comprehensive review of evaluation methodologies for LLM-based agents, marking a significant paradigm shift in autonomous AI systems. These agents, capable of planning, reasoning, tool usage, and maintaining memory in dynamic environments, require robust evaluation frameworks to gauge their performance accurately. The paper systematically dissects evaluation benchmarks and methodologies across four critical dimensions: assessing fundamental capabilities such as planning, self-reflection, tool utilization, and memory; domain-specific benchmarks for areas like web interactions, software engineering, scientific inquiry, and conversational tasks; evaluation standards for generalist agents; and overarching frameworks that integrate these diverse facets. Notably, the survey highlights emerging trends toward more realistic, challenging evaluations with continuously updated benchmarks while identifying key gaps, especially regarding cost-efficiency, safety, robustness, and scalability that must be addressed. This work thus not only maps the rapidly evolving landscape of agent evaluation and sets forth a clear research agenda for advancing future methodologies.
@@ II-F Our Survey
Our study makes a series of novel contributions compared to related works as presented in Table. I. We analyze the top 27 LLM models released between 2023 and 2025, including models such as Mistral AI Small 3 24B, DeepSeek-R1, Search-o1, QwQ-32B, and phi-4, providing detailed insights into their architectures, performance, and innovations. In addition, we offer a thorough overview of training methodologies by categorizing approaches into general training techniques, mixture-of-experts (MoE) innovations, retrieval-augmented generation (RAG), chain-of-thought and self-improvement methods, as well as test-time compute scaling, distillation, and reinforcement learning (RL) strategies.
@@
[…]
@@
[…]
@@ III Top LLM models
Recent advances in LLMs during 2023–2025 have demonstrated remarkable progress in architecture and training methodologies. For example, UI-TARS [qin2025ui] and Tülu 3 [lambert2025tulu3pushingfrontiers] exemplify novel approaches in multimodal processing and transparent post-training pipelines, respectively, that have enabled significant improvements in reasoning, code generation, and domain-specific tasks. These developments leverage innovations such as reinforcement learning with verifiable rewards, retrieval-augmented generation, and large-scale fine-tuning, addressing critical scalability, efficiency, and task adaptability challenges. The concentrated efforts during this period underscore a broader trend toward enhancing performance through high-quality synthetic data and advanced optimization techniques.
In parallel, models such as HuatuoGPT-o1 [chen2024huatuogpt] and Search-o1 [li2025search] highlight a focused commitment to refining reasoning capabilities in specialized domains, including medicine and long-step problem-solving. In addition, breakthroughs in vision language integration and context management have been achieved by systems such as Qwen2.5 VL [qwen2.5_vl] and InternLM2 [cai2024internlm2], further broadening the scope of applications for LLM. These contributions, emerging in a concentrated timeframe between 2024 and 2025, reinforce the rapid pace of innovation in the field and provide a robust foundation for subsequent research and development in artificial intelligence.
@@ III-A UI-TARS model
UI-TARS [qin2025ui] is a novel native GUI agent model that exclusively utilizes screenshot inputs to execute human-like interactions such as mouse and keyboard operations. Unlike conventional frameworks that rely on heavily wrapped commercial models and meticulously designed prompts, UI-TARS has been developed as an end-to-end system that surpasses these sophisticated performance approaches. Experimental evaluations across more than ten GUI agent benchmarks assessing perception, grounding, and task execution demonstrate its state-of-the-art performance. For example, on the OSWorld benchmark, UI-TARS achieves scores of 24.6 and 22.7 for 50-step and 15-step evaluations, respectively, outperforming comparable models like Claude, while on AndroidWorld, it scores 46.6 compared to GPT-4’s 34.5. The system integrates several key innovations: it employs enhanced perception through a large-scale dataset of GUI screenshots for context-aware understanding; it utilizes unified action modeling to standardize interactions across platforms; it incorporates System-2 reasoning to enable deliberate multi-step decision-making via task decomposition, reflection, and milestone recognition; and it leverages iterative training with reflective online traces to continuously refine its performance with minimal human intervention. Available in 2B, 7B, and 72B parameter configurations, UI-TARS exemplifies a significant step forward in building versatile, reasoning-capable GUI agents that can perform complex interactions across diverse computing environments.
@@ III-B Tülu 3 405B model
The recent study on Tulu 3 [lambert2025tulu3pushingfrontiers] presents a significant advancement in language model post-training, offering a fully transparent and reproducible framework for refining and enhancing LLMs. Unlike proprietary post-training pipelines, which often lack transparency regarding training data and methodologies, Tulu 3 provides a comprehensive open-source alternative, sharing the model weights and the complete post-training recipe, including datasets, training code, infrastructure, and evaluation tools. Built upon Llama 3.1 base models, Tulu 3 surpasses existing instruct-tuned models such as Llama 3.1-Instruct, Qwen 2.5, Mistral, and even proprietary models like GPT-4o-mini and Claude 3.5-Haiku in multiple benchmarks. The training methodology incorporates Supervised Fine-tuning (SFT), Direct Preference Optimization (DPO), and an innovative technique termed Reinforcement Learning with Verifiable Rewards (RLVR), which ensures reliable and measurable improvements in model behavior. To further enhance post-training effectiveness, the authors introduce a multi-task evaluation framework, which includes development and unseen benchmarks, standardized performance metrics, and rigorous decontamination of existing open datasets.